OpenStudy (anonymous):

Evaluate. (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (anonymous):

(5 • 8 – 12)(36 ÷ 9 • 4) =(40-12)(4*4) =28*16 =448